# FedALT: Federated Fine-Tuning through Adaptive Local Training with Rest-of-the-World LoRA

### v1

@@ 4 Experiments
Table 2: Performance comparison with baseline methods in the second federated dataset setting. The average column represents the overall performance across these tasks. See details in Section 4.1.
4.1 Experiments Results
Overall Performance Comparisons. We compare the performance of the proposed method, FedALT, with other baseline methods based on each client’s performance on its specific test dataset, aligned with its training dataset. As shown in Table 1, which reports the results for the first federated setting, our proposed method achieves the best overall performance compared to all baselines, with the average score exceeding that of the best-performing baseline, FDLoRA, by 2.13. FedALT consistently delivers superior performance across most tasks, outperforming the baselines by at least 3.27 on the Natural Language Inference task and 3.50 on the Reading Comprehension task. We find that on the Struct-to-Text, the Local Only method achieves the highest performance. This is likely due to the unique nature of this task, making it difficult to benefit from shared global knowledge. Furthermore, we observe that FedIT underperforms compared to Local Only on four client tasks, indicating the presence of harmful interference during aggregation. Personalized federated LoRA fine-tuning methods (i.e., FedDPA, PF2LoRA, FDLoRA) also fall short compared to FedALT, likely because they fail to effectively isolate harmful interference during the server aggregation step. Despite their designs for local/global LoRA interaction, these methods update LoRA modules using aggregated global information, which may carry conflicting information from other clients.
To further validate the effectiveness of FedALT, Table 2 presents the results for the second federated setting, involving a different mix of NLP tasks, clients, and task types. The results align with those in the first setting, with FedALT outperforming all baseline methods. Notably, FedALT achieves an average score of 70.48, surpassing the baseline method by at least 2.88. This consistent performance across diverse settings highlights the robustness and adaptability of FedALT, demonstrating its ability to handle heterogeneous data distributions effectively.
Additionally, in Table 3, we report the number of trainable parameters and the parameters required for inference for both the baseline methods and our proposed method. Due to space limitations, we present FedIT as an example, as the other federated learning-based methods have identical trainable and inference parameters. All methods exhibit a similar number of trainable parameters, except for FFA-LoRA, which has the fewest trainable parameters since it only updates the LoRA $\mathbf{B}$ module. However, despite its reduced parameter count, FFA-LoRA is significantly outperformed by our proposed method in overall performance. Regarding the parameters required for inference, our method involves an additional computation for the RoTW LoRA. However, this parameter overhead is negligible compared to the size of the large pre-trained LLM.
Methods Trainable Parameters Additional Inference Workload
FedIT 0.0622% 0.0622%
FFA-LoRA 0.0311% 0.0311%
FedDPA 0.0622% 0.1245%
FedALT 0.0661% 0.1283%
Table 3: Comparison of trainable and inference parameters. Note that, except for FFA-LoRA and FedDPA, all other baselines share the same parameter count as FedIT. Our method employs a comparable number of LoRA parameters to other baselines, which is negligible compared to the pre-trained LLM. See details in Section 4.1.
Impact of Decoupling LoRA Training. In our proposed method, only the Individual LoRA modules are updated locally during client fine-tuning, while the RoTW LoRA remains frozen and is updated solely through server-side aggregation. This design aims to mitigate harmful cross-client interference by preventing local updates to the shared global component. To validate the effectiveness of this decoupled training approach, we compare FedALT against two alternative methods. The first comparison method, referred to as RoTW-Update (rank = 4), fine-tunes both the Individual LoRA and RoTW LoRA locally, with each LoRA having a rank of 4 to maintain an equivalent total number of trainable parameters. The second comparison method, RoTW-Update (rank = 8), also fine-tunes both LoRA modules locally but assigns a rank of 8 to each, resulting in a higher total parameter count. The results, shown in Table 4, reveal that while the second method outperforms the first due to its larger parameter count, both alternative methods underperform compared to our proposed approach. Notably, even with more trainable parameters, the second method fails to match the performance of FedALT.
Methods Same Trainable Param. Same Inference Param. Avg. Performance
RoTW-Update (rank = 4) $\checkmark$ $\times$ 63.59
RoTW-Update (rank = 8) $\times$ $\checkmark$ 65.75
FedALT 67.55
Table 4: Impact of decoupling LoRA training. Updating the RoTW LoRA component locally leads to harmful cross-client interference. See details in Section 4.1.
Figure 3: Impact of Dynamic Weighting via MOE. The proposed dynamic personalized mixer mechanism surpasses both fixed-weight approaches and shared averaged mixers across clients. See Section 4.1 for details.
Impact of Dynamic Weighting via MOE. To enable input-specific flexibility, our proposed method adopts a MOE mechanism to dynamically adjust the contributions of the Individual LoRA and RoTW LoRA. We conduct experiments to evaluate the efficiency of this dynamic weighting by comparing it to two fixed weighting strategies. In the first fixed strategy, we set the contribution weight to a fixed value of $\alpha=0.5$, meaning that the Individual LoRA and RoTW LoRA contribute equally. In the second fixed strategy, we set the contribution weight to $\alpha=1/K$, where $K$ is the number of clients. This configuration effectively reduces the method to using the averaged LoRA, relying solely on each client’s last-round LoRA modules for the next round’s initialization. The results in Figure 3 validate that our proposed dynamic weighting mechanism outperforms both fixed strategies, demonstrating its ability to enhance performance. Additionally, our dynamic weighting is implemented through the personal $\mathbf{G}_{k}$ mixer, which is maintained locally by each client and is not aggregated during the training process. To further investigate this design choice, we conducted experiments where $\mathbf{G}_{k}$ is averaged across clients. The results confirm that retaining a personalized $\mathbf{G}_{k}$ for each client yields optimal performance, emphasizing the importance of input-specific and client-specific adjustments.
5 Related Work
5.1 Parameter-Efficient Fine-Tuning
As the size of LLMs continues to grow, full parameter fine-tuning has become increasingly computationally and storage-intensive (Hadi et al., 2023). To address this issue, Parameter-Efficient Fine-Tuning (PEFT) methods (Ding et al., 2023; Fu et al., 2023; Han et al., 2024; Liu et al., 2022; Li & Liang, 2021; Lester et al., 2021) have been developed to significantly reduce the number of trainable parameters. PEFT techniques introduce a limited set of additional trainable parameters to enhance model performance while keeping the majority of pre-trained parameters frozen. Among all existing PEFT methods, the most prominent approach is LoRA (Hu et al., 2021), which employs low-rank matrices to approximate the pre-trained weight matrix, updating only the low-rank components.
LoRA has gained significant attention recently, becoming a standard method for adapting LLMs like Llama under limited computational resources (Lermen et al., 2023). Several works have been proposed to improve LoRA (Tian et al., 2024; Lermen et al., 2023; Wang et al., 2023; Sheng et al., 2023; Liu et al., 2023). For instance, a recent development, Hydra-LoRA (Tian et al., 2024), employs an asymmetric structure with a shared matrix for all samples and distinct matrices for each intrinsic component, thereby enhancing domain adaptation. While these advancements primarily focus on centralized learning settings, our work extends the application of PEFT methods to federated LLM fine-tuning, addressing the unique challenges posed by distributed and heterogeneous client data.
5.2 Federated Learning
The foundational work in federated learning, FedAvg (McMahan et al., 2017), addresses privacy and communication efficiency by aggregating local model parameters to train a shared global model. Since then, numerous studies (Wang et al., 2024a, c; Yang et al., 2024b; Bian et al., 2024a) have focused on tackling various challenges within FL settings. One of the key challenges is data heterogeneity, which makes it difficult to train a single shared global model that performs well across all clients. To address this, Personalized Federated Learning(PFL) (Tan et al., 2022; Deng et al., 2020; Collins et al., 2021; Fallah et al., 2020) has emerged as an approach to adapt the global model to the specific needs of each client in the federation. Most PFL research has focused on addressing statistical differences in data distributions (e.g., label distribution skew) and has been applied primarily to smaller, standard models. In contrast, our work addresses the unique challenges posed by heterogeneous clients with diverse NLP tasks by leveraging LLM fine-tuning within the FL framework.
5.3 Federated Fine-Tuning
Several studies (Bai et al., 2024; Cho et al., 2024; Wang et al., 2024d; Kuang et al., 2024; Wu et al., 2024; Bian et al., 2024b) have explored federated fine-tuning approaches. (Kuang et al., 2024) proposed federated fine-tuning with all parameters updated, while (Sun et al., 2022) introduced federated fine-tuning with PEFT using prefix-tuning. (Zhang et al., 2024) was the first study to apply LoRA in a federated context. (Bai et al., 2024; Cho et al., 2024) focused on federated fine-tuning where clients have different LoRA ranks, while (Wang et al., 2024d) proposed strategies for LoRA initialization to achieve better performance. Another studies (Bian et al., 2024b; Sun et al., 2024) addressed server aggregation bias in LoRA-based federated fine-tuning.
The most relevant works to ours focus on personalized federated fine-tuning with LoRA (Anonymous, 2025; Yang et al., 2024a; Qi et al., 2024). FedDPA (Yang et al., 2024a) introduced global and local LoRA components but sufferd from interference during global LoRA training and struggled to balance the two phases effectively. PF2LoRA (Anonymous, 2025) employed a bilevel framework to combine a shared adapter with client-specific adapters but faced interference issues during the aggregation of the shared adapter. FDLoRA (Qi et al., 2024) utilized personalized LoRA modules to initialize federated training, combining them with a global LoRA module via adaptive fusion; however, it relied on server-side datasets and was prone to interference during global LoRA aggregation, limiting its effectiveness in heterogeneous settings.
6 Conclusion
In this work, We proposed FedALT, a personalized federated LoRA fine-tuning framework that departs from the FedAvg paradigm by prioritizing local adaptation while integrating global knowledge via a Rest-of-the-World (RoTW) LoRA component. An adaptive mixer dynamically balances local and global information, mitigating cross-client interference. Experiments across NLP tasks demonstrate FedALT’s superiority in heterogeneous settings with minimal overhead. Future work will focus on a theoretical analysis of FedALT’s convergence.
Appendix A Datasets
In this paper, we primarily follow the setup of (Yang et al., 2024a) to construct federated datasets for heterogeneous client settings using the Flan datasets (Chung et al., 2024). The Flan collection consists of over 12 NLP tasks, each containing multiple datasets. These tasks, generated from diverse contextual factors, inherently experience various complex distribution shifts. We consider two federated dataset configurations. For the first configuration, we assume an extreme client heterogeneity setting, where 8 clients are assigned completely different NLP tasks (a total of 8 distinct tasks are considered). For the second configuration, we consider a relatively moderate setting, where some clients share the same NLP task but with data originating from different datasets. In this setting, 8 clients with 6 distinct NLP tasks are included. To simulate a realistic FL environment, where each client has limited access to large datasets, we employ a downsampling strategy, assigning each client 600 training samples and 300 testing samples. The specific tasks and datasets for each federated dataset configuration are detailed in Tables 5 and 6. Note that while there is some overlap in the datasets used for federated setting 1 and federated setting 2, the samples within each configuration are randomly selected, ensuring independence between the two settings.
The original NLP tasks in Flan exhibit diverse formats, making it challenging to standardize processing for LLMs and to select a consistent evaluation metric. Even within a single dataset (e.g., Sentiment140), samples can vary significantly in format, as illustrated in Tables 7 and 8. To facilitate uniform processing and evaluation, we adopt a simple prompt template following (Yang et al., 2024a), as shown in Table 9. ROUGE-1 is chosen as the evaluation metric to assess performance across all tasks.
Task Dataset Description
Text Classification ag_news_subset News classification into categories.
Natural Language Inference snli Determine relationships between sentence pairs.
Reading Comprehension openbookqa Open-book question answering.
Paraphrase Detection glue_mrpc Identify semantic equivalence in sentences.
Commonsense Reasoning story_cloze Choose the correct story ending.
Struct to Text common_gen Generate text from concepts.
Sentiment Analysis sentiment140 Sentiment analysis of tweets.
Coreference Resolution definite_pronoun_resolution Resolve ambiguous pronouns.
Table 5: Tasks and Datasets for Federated Dataset 1
Task Dataset Description
Struct to Text 1 common_gen Generate text from concepts.
Struct to Text 2 dart Generate structured text descriptions.
Coreference Resolution definite_pronoun_resolution Resolve ambiguous pronouns.
Natural Language Inference 1 snli Determine relationships between sentence pairs.
Natural Language Inference 2 qnli Answer questions based on context.
Paraphrase Detection glue_mrpc Identify semantic equivalence in sentences.
Text Classification ag_news_subset Classify news into categories.
Sentiment Analysis sst2 Analyze sentiment of text.
Table 6: Tasks and Datasets for Federated Dataset 2
Data Examples
Instruction: @celinieee The crafts thingy is already done. Describe the sentiment embodied by this tweet. OPTIONS: - negative - positive.
Output: negative
Dataset: sentiment140
Instruction: Write a positive tweet.
Output: @MandinaM Very well indeed thank you.. Yourself? Just about to start work.. 9 hours to go..
Dataset: sentiment140
Table 7: Examples of Data in Sentiment Analysis Task.
Data Examples
Instruction: What are the most important words in the following sentence: pig in a pen on a pig farm
Output: farm, pen, pig
Dataset: common_gen
Instruction: Generate a sentence about this data: Wildwood eatType coffee shop; Wildwood priceRange moderate; Wildwood customer rating 1 out of 5; Wildwood near Ranch
Output: The Wildwood coffee shop is an average-priced location that serves a light menu of food and alcohol. Located near the Ranch, customers give this location a 1 out of 5 stars.
Dataset: dart
Table 8: Examples of Data in Struct to Text Task.
Template
Prompt Input Instruction: {instruction} Response:
Table 9: Prompt Template.
Appendix B Training Implementations
We implement all methods using PyTorch and conduct all experiments on six NVIDIA A100 GPUs. For the federated methods, we set the local epochs to 5 and the global rounds to 20. For the Local Only baseline, the total number of training rounds is set to 50. The learning rate is set to $3\times 10^{-4}$, and the AdamW optimizer is utilized, following (Yang et al., 2024a).
LoRA modules are inserted into the $q$ and $v$ components of each self-attention layer in LLaMA2. The batch size is set to 8, and the rank of the trainable LoRA for each method is fixed at 8. Additionally, we set lora_alpha to 32 and lora_dropout to 0.05. Due to computational limitations, the pre-trained LLaMA2 model is loaded in 8-bit precision.
Appendix C Methods
Local Only. The Local Only method represents a setting where each client independently trains its own model without sharing any model updates with a server. This method serves as a baseline to assess the impact of federated learning methods on model performance.
FedIT (Zhang et al., 2024). FedIT integrates LoRA with the FedAvg framework, making it one of the earliest approaches to combine these techniques. However, under heterogeneous client data settings, FedIT suffers from harmful cross-client interference during the aggregation step, as highlighted in the motivational experiments. This interference often reduces its efficiency and, in some cases, results in worse performance than the Local Only method.
FFA-LoRA (Sun et al., 2024). FFA-LoRA aims to reduce communication costs by fixing the randomly initialized non-zero $\mathbf{A}$ matrices and fine-tuning only the zero-initialized $\mathbf{B}$ matrices. While this approach lowers communication overhead, it impairs the learning capacity of LoRA, as fixed $\mathbf{A}$ matrices limit the model’s ability to adapt to diverse client data, resulting in suboptimal performance.
FedSA (Guo et al., 2024). FedSA introduces an asymmetric approach to LoRA matrix updates in federated learning, where only the $\mathbf{A}$ matrices are shared with the server for aggregation, while the $\mathbf{B}$ matrices remain local to capture client-specific knowledge. However, even the client-specific $\mathbf{B}$ matrices may contain valuable knowledge that could benefit other clients. The failure to leverage this knowledge effectively results in reduced overall performance.
FedDPA (Yang et al., 2024a). FedDPA is a state-of-the-art personalized federated fine-tuning method that introduces two distinct LoRA components: global LoRA and local LoRA. The global LoRA is trained using FedAvg, while the local LoRA is independently fine-tuned on client-specific data. Despite its personalization capabilities, FedDPA suffers from interference during the global LoRA training stage. Additionally, its separation of local and global LoRA fine-tuning into distinct phases complicates the determination of optimal training durations for each phase.
[…]

### v2

@@ E.2 Per-Client Performance
Table 14: Performance on mild heterogeneous (Bloom-560M). Evaluation across tasks in Table 6.
Methods Bloom-560M Average
Struct 1 Struct 2 Coref. NLI 1 NLI 2 Para. Text Cls. Sentiment
Local 40.47 48.10 34.48 42.33 52.58 70.26 47.65 65.82 50.21
FedIT 35.38 41.20 38.86 39.23 49.16 71.67 51.75 65.00 49.03
FFA-LoRA 36.16 39.40 36.57 36.22 47.19 68.67 50.45 63.89 47.32
FedSA 35.98 44.71 38.68 47.68 55.92 70.54 58.02 68.84 52.55
FedDPA 38.56 42.54 39.23 46.69 49.21 69.00 62.35 68.92 52.06
PF2LoRA 39.65 42.29 40.36 45.85 53.07 70.40 57.43 66.67 51.97
FDLoRA 40.66 44.62 38.91 43.54 51.17 70.33 54.13 68.00 51.42
FedALT 41.10 48.64 39.80 48.14 56.63 71.27 64.84 68.67 54.89
Table 15: Performance on low heterogeneous (LLaMA2-7B). Evaluation across 8 clients split by task-dataset pairs. SA1–SA2 use sst2, SA3–SA4 use sentiment140, NLI1–NLI2 use qnli, and NLI3–NLI4 use snli.
Methods SA1 SA2 SA3 SA4 NLI1 NLI2 NLI3 NLI4 Average
Local 65.95 63.95 63.57 66.34 45.95 47.45 66.93 64.60 60.59
FedIT 67.25 66.82 65.85 64.79 46.53 49.48 68.53 68.65 62.24
FFA-LoRA 66.42 61.84 62.96 65.14 46.83 49.23 67.93 68.16 61.07
FedSA 71.85 64.31 65.45 69.37 46.31 48.24 67.03 70.38 62.87
FedDPA 67.39 64.74 65.58 63.98 46.89 48.63 69.77 69.47 62.06
PF2LoRA 71.73 67.83 65.45 70.18 47.19 47.68 69.41 69.03 63.56
FDLoRA 68.54 65.46 65.93 63.36 47.09 49.19 70.69 69.07 62.42
FedALT 73.05 67.86 65.91 72.82 47.29 50.01 70.42 69.88 64.66
Table 16: Performance on low heterogeneous (Bloom-560M). Each dataset is split across two clients. See Table 7 for dataset mapping.
Methods SA1 SA2 SA3 SA4 NLI1 NLI2 NLI3 NLI4 Average
Local 44.33 40.45 46.03 51.04 40.64 42.58 62.13 62.03 48.65
FedIT 46.89 45.16 50.20 54.69 42.49 46.26 64.19 67.29 52.15
FFA-LoRA 52.18 36.04 52.47 51.91 42.78 45.29 64.50 65.45 51.33
FedSA 53.55 47.16 51.63 50.47 44.96 45.92 66.36 64.18 53.03
FedDPA 52.61 47.27 52.03 53.14 43.26 45.56 65.68 65.38 53.12
PF2LoRA 46.88 43.00 55.24 55.30 39.42 43.39 64.14 61.39 51.10
FDLoRA 50.66 46.72 49.38 54.07 42.35 45.82 64.63 66.45 52.51
FedALT 54.06 49.18 53.34 55.78 42.94 45.51 67.57 67.49 54.48
Appendix F Convergence Analysis
Since our mixer depends on each sample input, existing convergence analysis frameworks cannot be directly applied. In this section, we provide an initial convergence analysis of FedALT under appropriate assumptions. We begin by defining the key parameters and establishing necessary conditions. While this analysis offers theoretical validation for our approach, we acknowledge that it represents a preliminary investigation into the convergence properties of FedALT. More comprehensive theoretical frameworks with potentially stronger guarantees and less restrictive assumptions remain an important direction for future research. Nevertheless, our initial analysis provides valuable insights into the convergence behavior and offers a theoretical foundation that complements our empirical findings.
F.1 Notation and Setup
Let the trainable parameters for client $k$ at round $t$ be denoted by
$Z_{k}^{t}=\{\mathbf{A}_{k}^{L},\mathbf{B}_{k}^{L},\mathbf{G}_{k}\}$ (5)
where $\mathbf{A}_{k}^{L}$ and $\mathbf{B}_{k}^{L}$ represent the Individual LoRA modules and $\mathbf{G}_{k}$ denotes the adaptive mixer.
The Rest-of-World (RoW) LoRA for client $k$ at round $t$ is computed as:
$\mathbf{A}_{k}^{R}=\frac{1}{K-1}\sum_{m\neq k}\mathbf{A}_{m}^{L},\quad\mathbf{B}_{k}^{R}=\frac{1}{K-1}\sum_{m\neq k}\mathbf{B}_{m}^{L}$ (6)
We denote the loss function for client $k$ as $L_{k}(Z_{k}|\mathbf{A}_{k}^{R},\mathbf{B}_{k}^{R})$, which measures the performance on client $k$’s local dataset given its trainable parameters and the fixed RoW LoRA components.
F.2 Assumptions
Assumption F.1 (Smoothness).
For each client $k$, the loss function $L_{k}$ is $\mu$-smooth with respect to $Z_{k}$, i.e., for any $Z_{k}^{1},Z_{k}^{2}$:
$\|\nabla L_{k}(Z_{k}^{1})-\nabla L_{k}(Z_{k}^{2})\|\leq\mu\|Z_{k}^{1}-Z_{k}^{2}\|$ (7)
Assumption F.2 (Strong Convexity).
For each client $k$, the loss function $L_{k}$ is $\lambda$-strongly convex with respect to $Z_{k}$, i.e., for any $Z_{k}^{1},Z_{k}^{2}$:
$L_{k}(Z_{k}^{2})\geq L_{k}(Z_{k}^{1})+\langle\nabla L_{k}(Z_{k}^{1}),Z_{k}^{2}-Z_{k}^{1}\rangle+\frac{\lambda}{2}\|Z_{k}^{2}-Z_{k}^{1}\|^{2}$ (8)
Assumption F.3 (Bounded RoW Sensitivity).
The optimal parameters for client $k$ are sensitive to changes in the RoW parameters with a bounded Lipschitz constant. Specifically, for any two sets of RoW parameters $(\mathbf{A}_{k}^{R,1},\mathbf{B}_{k}^{R,1})$ and $(\mathbf{A}_{k}^{R,2},\mathbf{B}_{k}^{R,2})$:
$\|Z_{k}^{*}(\mathbf{A}_{k}^{R,1},\mathbf{B}_{k}^{R,1})-Z_{k}^{*}(\mathbf{A}_{k}^{R,2},\mathbf{B}_{k}^{R,2})\|\leq\beta\|(\mathbf{A}_{k}^{R,1},\mathbf{B}_{k}^{R,1})-(\mathbf{A}_{k}^{R,2},\mathbf{B}_{k}^{R,2})\|$ (9)
where $Z_{k}^{*}(\mathbf{A}_{k}^{R},\mathbf{B}_{k}^{R})=\arg\min_{Z_{k}}L_{k}(Z_{k}|\mathbf{A}_{k}^{R},\mathbf{B}_{k}^{R})$ represents the optimal parameters for client $k$ given fixed RoW parameters.
The first two assumptions are standard and widely used in FL proof. The third assumption holds because the RoW LoRA component remains frozen during local training, providing a stable reference point that limits how much client parameters can drift between rounds based on changes in other clients’ parameters.
F.3 Convergence Theorem
Theorem F.4 (Convergence of FedALT).
With the assumptions, we can derive:
$\|Z_{k}^{t+1}-Z_{k}^{t}\|\leq\frac{2\epsilon}{1-\beta}+\beta^{t}\max_{k}\|Z_{k}^{1}-Z_{k}^{0}\|$
If $\beta<1$ and the local training at each round converges to a neighborhood of the optimal solution, then the sequence of trainable parameters $\{Z_{k}^{t}\}$ generated by FedALT converges to a stable point for each client $k$.
Proof.
We analyze the convergence in two steps:
  1. 1.
    Local convergence within each round
  2. 2.
    Global convergence across rounds
Step 1: Local convergence within each round
At round $t$, client $k$ optimizes its parameters $Z_{k}$ with fixed RoW parameters $(\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})$. Let $Z_{k}^{t,s}$ represent the parameters after $s$ steps of local training. We now analyze the convergence of this local optimization process using Assumptions 1 and 2.
The gradient descent update rule at step $s$ is given by:
$Z_{k}^{t,s+1}=Z_{k}^{t,s}-\eta\nabla L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})$ (10)
From Assumption 1 (smoothness), we can have:
$L_{k}(Z_{k}^{t,s+1}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})$ $\leq L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})+\langle\nabla L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t}),Z_{k}^{t,s+1}-Z_{k}^{t,s}\rangle$
$\quad+\frac{\mu}{2}\|Z_{k}^{t,s+1}-Z_{k}^{t,s}\|^{2}$ (11)
Substituting the update rule:
$L_{k}(Z_{k}^{t,s+1}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})$ $\leq L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})$
$\quad+\langle\nabla L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t}),-\eta\nabla L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})\rangle$
$\quad+\frac{\mu}{2}\|-\eta\nabla L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})\|^{2}$ (12)
$=L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})-\eta\|\nabla L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})\|^{2}$
$\quad+\frac{\mu\eta^{2}}{2}\|\nabla L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})\|^{2}$ (13)
$=L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})-\eta(1-\frac{\mu\eta}{2})\|\nabla L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})\|^{2}$ (14)
Let $Z_{k}^{*}=Z_{k}^{*}(\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})$ denote the optimal parameters that minimize the loss function for client $k$ given the fixed RoW parameters.
From Assumption 2 (strong convexity), we have:
$L_{k}(Z_{k}^{*}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})$
$\geq$ $L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})+\langle\nabla L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t}),Z_{k}^{*}-Z_{k}^{t,s}\rangle+\frac{\lambda}{2}\|Z_{k}^{*}-Z_{k}^{t,s}\|^{2}$ (15)
Rearranging:
$\langle\nabla L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t}),Z_{k}^{t,s}-Z_{k}^{*}\rangle$
$\geq$ $L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})-L_{k}(Z_{k}^{*}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})+\frac{\lambda}{2}\|Z_{k}^{*}-Z_{k}^{t,s}\|^{2}$ (16)
By Cauchy-Schwarz inequality:
$\|\nabla L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})\|\cdot\|Z_{k}^{t,s}-Z_{k}^{*}\|\geq\langle\nabla L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t}),Z_{k}^{t,s}-Z_{k}^{*}\rangle$ (17)
Combining with the previous inequality:
$\|\nabla L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})\|\cdot\|Z_{k}^{t,s}-Z_{k}^{*}\|$
$\geq$ $L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})-L_{k}(Z_{k}^{*}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})+\frac{\lambda}{2}\|Z_{k}^{*}-Z_{k}^{t,s}\|^{2}$ (18)
From the properties of strongly convex functions, we can derive:
$\|\nabla L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})\|^{2}\geq 2\lambda(L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})-L_{k}(Z_{k}^{*}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t}))$ (19)
Substituting this into our earlier expression for the loss after one update:
$L_{k}(Z_{k}^{t,s+1}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})-L_{k}(Z_{k}^{*}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})$
$\leq L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})-L_{k}(Z_{k}^{*}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})-\eta(1-\frac{\mu\eta}{2})\|\nabla L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})\|^{2}$ (20)
$\leq L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})-L_{k}(Z_{k}^{*}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})$
$-2\eta(1-\frac{\mu\eta}{2})\lambda(L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})-L_{k}(Z_{k}^{*}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t}))$ (21)
$=(1-2\eta\lambda(1-\frac{\mu\eta}{2}))(L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})-L_{k}(Z_{k}^{*}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})),$ (22)
where the learning rate $\eta<\frac{2}{\mu}$.
From Assumption 2 (strong convexity), we have:
$\frac{\lambda}{2}\|Z_{k}^{t,s}-Z_{k}^{*}\|^{2}\leq L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})-L_{k}(Z_{k}^{*}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})$ (23)
And from Assumption 1 (smoothness):
$L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})-L_{k}(Z_{k}^{*}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})\leq\frac{\mu}{2}\|Z_{k}^{t,s}-Z_{k}^{*}\|^{2}$ (24)
Combining these inequalities with our convergence rate:
$\frac{\lambda}{2}\|Z_{k}^{t,s+1}-Z_{k}^{*}\|^{2}$ $\leq L_{k}(Z_{k}^{t,s+1}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})-L_{k}(Z_{k}^{*}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})$ (25)
$\leq(1-2\eta\lambda(1-\frac{\mu\eta}{2}))(L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})-L_{k}(Z_{k}^{*}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t}))$ (26)
$\leq(1-2\eta\lambda(1-\frac{\mu\eta}{2}))\frac{\mu}{2}\|Z_{k}^{t,s}-Z_{k}^{*}\|^{2}$ (27)
Simplifying:
$\|Z_{k}^{t,s+1}-Z_{k}^{*}\|^{2}\leq(1-2\eta\lambda(1-\frac{\mu\eta}{2}))\frac{\mu}{\lambda}\|Z_{k}^{t,s}-Z_{k}^{*}\|^{2}$ (28)
Denote $\rho=(1-2\eta\lambda(1-\frac{\mu\eta}{2}))\frac{\mu}{\lambda}$, when we set the learning rate be $\frac{1}{\mu}$, we can make $\rho$ in the range of 0 to 1.
By applying this inequality recursively for $s$ steps:
$\|Z_{k}^{t,s}-Z_{k}^{*}\|^{2}\leq\rho^{s}\|Z_{k}^{t,0}-Z_{k}^{*}\|^{2}$ (29)
Taking the square root:
$\|Z_{k}^{t,s}-Z_{k}^{*}(\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})\|\leq\rho^{s/2}\|Z_{k}^{t,0}-Z_{k}^{*}(\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})\|$ (30)
Since $\rho<1$, we have $\lim_{s\to\infty}\rho^{s}=0$. Therefore, with a sufficient number of local training steps $S$, we can ensure:
$\|Z_{k}^{t+1}-Z_{k}^{*}(\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})\|\leq\epsilon$ (31)
[…]